Lupo Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-hours. The company based its predetermined overhead rate for the current year on the following data:

Total machine-hours 33,000
Total fixed manufacturing overhead cost $ 660,000
Variable manufacturing overhead per machine-hour $ 6

Recently, Job T687 was completed with the following characteristics:

Number of units in the job 10
Total machine-hours 40
Direct materials $ 715
Direct labor cost $ 1,430

The total job cost for Job T687 is closest to: (Round your intermediate calculations to 2 decimal places.)

Garrison 16e Rechecks 2017-06-28

Multiple Choice

  • $2,470

  • $2,145

  • $1,755

  • $3,185
Answer #1

Predetermined Overhead rate

= Variable overhead rate + Fixed overhead rate

= 6 + (660,000/33000)

= 26 per machine hour

Total manufacturing cost = Direct Materials + Direct Labor + Overhead applied

= 715 + 1430 + (40*26)

= 715 + 1430 + 1040

= $3,185